OpenStudy (anonymous):

Solve for x. x divided by seven minus eight equals negative seventeen x equals negative sixty-three x equals negative sixteen x equals sixteen x equals sixty-three

OpenStudy (jadeishere):

x/7 - 8 = -17 x/7 = 9 x = 9 * 7 x =63

OpenStudy (jadeishere):

-63 sorry

OpenStudy (anonymous):

Xavier's dining room table was unstable because one leg was shorter than the others. He placed a 3/8 inch thick piece of wood under the short leg, and it became even with the other legs, which are 19 1/4 inches long. How many inches long was the short leg before Xavier added the wood? 19 5/8 inches 18 3/4 inches 19/ 7/8 inches 18 7/8 inches